# On the asymptotics of orthogonal polynomials on multiple intervals with non-analytic weights

### v2

@@ Abstract.
We consider the asymptotics of orthogonal polynomials for measures that are differentiable, but not necessarily analytic, multiplicative perturbations of Jacobi-like measures supported on disjoint intervals. We analyze the Fokas–Its–Kitaev Riemann–Hilbert problem using the Deift–Zhou method of nonlinear steepest descent and its $\overline{\partial}$ extension due to Miller and McLaughlin. Our results extend that of Yattselev in the case of Chebyshev-like measures with error bounds that are slightly improved while less regular perturbations are admissible. For the general Jacobi-like case, we present, what appears to be the first result for asymptotics when the perturbation of the measure is only assumed to be twice differentiable, with bounded third derivative.
@@ 1. Introduction
(1.1) $\mu(\mathop{}\!\mathrm{d}x)=\sum_{j=1}^{g+1}h_{j}(x)\mathbbm{1}_{[a_{j},b_{j}]}(x)(b_{j}-x)^{\alpha_{j}}(x-a_{j})^{\beta_{j}}\mathop{}\!\mathrm{d}x+\sum_{j=1}^{P}w_{j}\delta_{x_{j}}=\rho(x)\!\mathop{}\!\mathrm{d}x+\sum_{j=1}^{P}w_{j}\delta_{c_{j}},$
where $a_{j}<b_{j}<a_{j+1}$ and $\{c_{1},\ldots,c_{p}\}$ is a subset of the complement of $\bigcup_{j}[a_{j},b_{j}]$. We assume that $\alpha_{j}>-1,\beta_{j}>-1,w_{j}>0$ for each $j$ and $h_{j}(x)>0$ on $[a_{j},b_{j}]$. We consider the problem of determining the strong asymptotics of the (monic) orthogonal polynomials, denoted by $\pi_{n}(;\mu)$, with respect to $\mu$ as $n\to\infty$ when $h_{j}$ is not assumed to be analytic.
@@
The Riemann–Hilbert approach to determining strong asymptotics was pioneered in [22, 12, 9] for weights on the real axis and in [24] for the interval $[-1,1]$. Crucially, these works used that weight function $\rho$ in $\mu(\mathop{}\!\mathrm{d}x)=\rho(x)\mathop{}\!\mathrm{d}x$ had an appropriate analytic continuation to a neighborhood of the support of $\mu$. It is worth noting here that the approach in [25], for example, does not require analyticity.
The Riemann–Hilbert approach was extended to non-analyic weights on the unit circle and the real line in [26, 27]. Then, in [6], the approach was used to obtain asymptotics for orthogonal polynomials on $[-1,1]$ for smooth, but not analytic, perturbations of the Jacobi weight $(1-x)^{\beta}(1+x)^{\alpha}$. Continuing, in [37] the asymptotics were refined in the case where $\alpha,\beta\in\{-1/2,1/2\}$. This refinement is important because if $\rho(x)=h(x)(1-x)^{\beta}(1+x)^{\alpha}$, $\alpha,\beta\in\{-1/2,1/2\}$ and $h$ is analytic then one knows the error terms in the asymptotic expansions are exponentially small with respect to the degree of the polynomial [23]. The work of [37] uses a Riemann–Hilbert analysis to show that as the smoothness of $h$ increases, the order of the error term decreases accordingly. As noted in [37], this can also be seen to be a consequence of theorems concerning orthogonal polynomials on the unit circle, using the canonical (Joukowski) mapping of the unit circle to the unit interval $[-1,1]$.
@@
For general, $\alpha_{j},\beta_{j}$, we assume that $h_{j}\in C^{k,1}([a_{j},b_{j}])$, $k\geq 2$, with some further conditions on $k$ depending on $\alpha_{j},\beta_{j}$ (see Theorem 2.6). We also utilize the Sobolev spaces $W^{k,p}([a_{j},b_{j}])$ consisting of functions that have $k$ weak derivatives that are all in $L^{p}([a_{j},b_{j}])$ with respect to standard Lebesgue measure. In the special case where $\alpha_{j},\beta_{j}\in\{1/2,-1/2\}$ we allow $h_{j}\in W^{k,q}([a_{j},b_{j}])$, $k\geq 1$, $q>4$. Below, we reserve the index $k$ to refer to this assumed smoothness class.
As noted above, [37] considered the case of a single interval with $\alpha_{1},\beta_{1}\in\{1/2,-1/2\}$. It was shown that the error term is, in particular, $\mathrm{O}\left(\frac{\log n}{n^{k+\alpha}}\right)$ if $h_{1}\in C^{k,\alpha}([a_{1},b_{1}])$ and $k\geq 3$. Here we obtain an error bound of $\mathrm{O}\left(n^{-k-\alpha}\right)$ for $k\geq 1$, $q=\infty$, a modest improvement, while allowing $g>0$, $P>0$, see Theorem 3.1. There is a strong reason to believe this is still less than optimal, see Figures 4 and 5. These plots seem to indicate that this bound is suboptimal by a factor of $1/n$. It can be improved by allowing $p\approx 4$ and this room for improvement is discussed in Remark 3.2, but we feel it is still not sufficient to close this gap.
In the same vein as [37], the method described in this paper works by constructing close-to-analyic extensions of each function $h_{j}$ using polynomial approximation. Then, again borrowing directly from [37], we reuse an argument based on the Bernstein–Walsh inequality. Here it requires a bit more care in its implementation using a comparision, see Proposition A.1.
@@
The outline of the paper is as follows. In the remainder of this section, we introduce the Fokas–Its-Kitaev Riemann–Hilbert problem and fix more notation. In Section 2 we present the $\overline{\partial}$ deformations for the general Jacobi-type case $\alpha_{j},\beta_{j}>-1$. This leads to Theorem 2.6. In Section 3 we first present an improvement for the Chebyshev-like case $\alpha_{j},\beta_{j}\in\{-1/2,1/2\}$ in Theorem 3.1 and then discuss the addition of a finite number of point masses. Section 4 discusses the implied asymptotics for recurrence coefficients and uses this to present numerical experiments demonstrating that Theorem 3.1 is likely suboptimal. We also include five appendices that include techinical developments.
@@ 1.1. The Fokas–Its–Kitaev Riemann–Hilbert problem
(1.6) $\mathrm{Res}_{z=c_{j}}\mathbf{Y}_{n}(z;\mu)$ $=\lim_{z\to c_{j}}\mathbf{Y}_{n}(z;\mu)\begin{bmatrix}0&\frac{w_{j}}{2\pi\mathrm{i}}\\ 0&0\end{bmatrix},\quad j=1,2,\ldots,P.$
@@
(1.8) $\begin{split}\mathbf{Y}_{n}(z;\mu)&=\begin{cases}\mathrm{O}\begin{bmatrix}1&1+|z-b_{j}|^{\alpha_{j}}\\ 1&1+|z-b_{j}|^{\alpha_{j}}\end{bmatrix}&\alpha_{j}\neq 0,\vspace{4pt}\\ \mathrm{O}\begin{bmatrix}1&\log|z-b_{j}|\\ 1&\log|z-b_{j}|\end{bmatrix}&\alpha_{j}=0,\end{cases}\quad z\to b_{j},\\ \mathbf{Y}_{n}(z;\mu)&=\begin{cases}\mathrm{O}\begin{bmatrix}1&1+|z-a_{j}|^{\beta_{j}}\\ 1&1+|z-a_{j}|^{\beta_{j}}\end{bmatrix}&\beta_{j}\neq 0,\vspace{4pt}\\ \mathrm{O}\begin{bmatrix}1&\log|z-a_{j}|\\ 1&\log|z-a_{j}|\end{bmatrix}&\beta_{j}=0,\end{cases}\quad z\to a_{j}.\end{split}$
@@
will be analytic except for possibly $a_{j},b_{j}$ where isolated singularities could persist. Indeed, we check, as $z\to b_{j}$, if $\alpha_{j}\neq 0$
@@ 2. Dbar deformation
We begin the deformation procedure by first recalling the definition and properties of the so-called $g$-function, see (A.2). This function is also related to the exterior Green’s function with pole at $\infty$ [35]. Often this function is determined during the process of deformation, but we need it at the outset. Throughout what follows, we now suppose that $\epsilon^{\prime}>\epsilon>0$ are sufficiently small so that both $\varphi(z;a_{j}),\varphi(z;b_{j})$ from (A.4) are conformal on the balls $B_{\epsilon^{\prime}}(a_{j}),B_{\epsilon^{\prime}}(b_{j})$ respectively [^2] Here we use the notation $B_{\epsilon}(c)=\{z\in\mathbb{C}~{}:~{}|z-c|<\epsilon\}$, for all $j$. And that all the sets $B_{\epsilon^{\prime}}(a_{j}),B_{\epsilon^{\prime}}(b_{j})$, $j=1,2,\ldots,g+1$ are at a distance at least, say, $10\epsilon^{\prime}$ from one another. Next, define
@@
$\Gamma_{1}$$\Gamma_{3}$$\Gamma_{2}$$\sigma$00
@@
$\Sigma_{a_{j},3}$$\Sigma_{a_{j},1}$$\Sigma_{a_{j},2}$$a_{j}$
$\Sigma_{b_{j},1}$$\Sigma_{b_{j},3}$$\Sigma_{b_{j},2}$$b_{j}$
Figure 2. Mapping the contours $\Gamma_{j}$ using the local conformal mappings.
The contours in Figure 2 will become jump contours in a deformed hybrid $\overline{\partial}$ problem.
@@
In Appendix B, we develop a operator that extends a function off the real axis, while satisfying several convenient conditions. For $f_{j}(x)=1/h_{j}(x)$, consider $f_{j}^{\sharp}(x,y)$ with $\kappa=k$, $\tau=\epsilon$. The choice of $\theta_{1},\theta_{2}$ that appear in the definition of $f_{j}^{\sharp}$ will be discussed below.
@@
giving a contour $\Sigma_{j,-}$. And a straight line, lying within the upper-half plane, can be used to connect
@@
giving a contour $\Sigma_{j,+}$. This is depicted in Figure 3. The contours $\Sigma_{j,\pm}$ now enclose open regions $\Omega_{j,\pm}$, also depicted in Figure 3.
$\Sigma_{a_{j},3}$$\Sigma_{a_{j},1}$$\Sigma_{a_{j},2}$$a_{j}$$\Sigma_{b_{j},1}$$\Sigma_{b_{j},3}$$\Sigma_{b_{j},2}$$b_{j}$$\Sigma_{j,+}$$\Sigma_{j,-}$$\Omega_{j,+}$$\Omega_{j,-}$
@@
¯$\overline{\partial}=\frac{1}{2}(\partial_{x}+\mathrm{i}\partial_{y}).$
@@
    For $c=a_{j},b_{j}$, $|z-c|<\epsilon/2$, and $z\in\Sigma_{c,1}\cup\Sigma_{c,3}$, we have
    $f_{j}^{\sharp}(x,y)=f_{j}(z;c)=\sum_{\ell=0}^{k}\frac{f_{j}^{(\ell)}(c)}{\ell!}(z-c)^{\ell},$
@@
    For $(x,y)\in{\Omega_{j,+}\cup\Omega_{j,-}}$, |¯$|\overline{\partial}f_{j}^{\sharp}(x,y)|\leq C\|f_{j}\|_{C^{k,1}([a_{j},b_{j})}|y|^{k}$, for some $C>0$.

### v4

@@ Abstract.
We consider the asymptotics of orthogonal polynomials for measures that are differentiable, but not necessarily analytic, multiplicative perturbations of Jacobi-like measures supported on disjoint intervals. We analyze the Fokas–Its–Kitaev Riemann–Hilbert problem using the Deift–Zhou method of nonlinear steepest descent and its $\overline{\partial}$ extension due to Miller and McLaughlin. Our results extend that of Yattselev in the case of Chebyshev-like measures with error bounds that give similar rates while allowing less regular perturbations. For the general Jacobi-like case, we present, what appears to be the first result for asymptotics when the perturbation of the measure is only assumed to be differentiable with bounded second derivative.
@@ 1. Introduction
(1.1) $\mu(\mathop{}\!\mathrm{d}x)=\sum_{j=1}^{g+1}h_{j}(x)\mathbbm{1}_{[a_{j},b_{j}]}(x)(b_{j}-x)^{\alpha_{j}}(x-a_{j})^{\beta_{j}}\mathop{}\!\mathrm{d}x+\sum_{j=1}^{P}{r}_{j}\delta_{{c}_{j}}=\rho(x)\!\mathop{}\!\mathrm{d}x+\sum_{j=1}^{P}{r}_{j}\delta_{c_{j}},$
where $a_{j}<b_{j}<a_{j+1}$ and $\{c_{1},\ldots,c_{P}\}$ is a subset of the complement of $\bigcup_{j}[a_{j},b_{j}]$. We assume that $\alpha_{j}>-1,\beta_{j}>-1,{r}_{j}>0$ for each $j$ and $h_{j}(x)>0$ on $a_{j},b_{j}]$. We consider the problem of determining the strong asymptotics of the (monic) orthogonal polynomials, denoted by $\pi_{n}(;\mu)$, with respect to $\mu$ as $n\to\infty$ when $h_{j}$ is not assumed to be analytic. Polynomials orthogonal over disjoint intervals were originally considered by Akhiezer [1] and Akhiezer and Tomchuck [2]. These authors described the orthogonal polynomials and their associated weighted Cauchy integrals, two of the main objects of study here.
@@
The Riemann–Hilbert approach to determining strong asymptotics was pioneered in [25, 15, 12] for weights on the real axis and in [27] for the interval [1,1][-1,1]. Crucially, these works used that weight function $\rho$ in $\mu(\mathop{}\!\mathrm{d}x)=\rho(x)\mathop{}\!\mathrm{d}x$ had an appropriate analytic continuation to a neighborhood of the support of $\mu$. It is worth noting here that the approach in [28], for example, does not require analyticity. For a special choice of weight function, explicit formulae exist in terms of Riemann theta functions [11].
The Riemann–Hilbert approach was extended to non-analytic weights on the unit circle and the real line in [29, 30]. Then, in [8], the approach was used to obtain asymptotics for orthogonal polynomials on [1,1][-1,1] for smooth, but not analytic, perturbations of the Jacobi weight $(1-x)^{\beta}(1+x)^{\alpha}$. Continuing, in [40] the asymptotics were refined in the case where $\alpha,\beta\in\{-1/2,1/2\}$. This refinement is important because if $\rho(x)=h(x)(1-x)^{\beta}(1+x)^{\alpha}$, $\alpha,\beta\in\{-1/2,1/2\}$ and $h$ is analytic then one knows the error terms in the asymptotic expansions are exponentially small with respect to the degree of the polynomial [26]. The work of [40] uses a Riemann–Hilbert analysis to show that as the smoothness of $h$ increases, the order of the error term decreases accordingly. As noted in [40], this can also be seen to be a consequence of theorems concerning orthogonal polynomials on the unit circle, using the canonical (Joukowski) mapping of the unit circle to the unit interval [1,1][-1,1].
@@
For general, $\alpha_{j},\beta_{j}$, we assume that $h_{j}\in C^{k,1}([a_{j},b_{j}])$, $k\geq{1}$ and $k>2\max_{j}\max\{|\alpha_{j}|,|\beta_{j}|\}-\frac{1}{2}$ (see Theorem 2.6). We also utilize the Sobolev spaces $W^{k,p}([a_{j},b_{j}])$ consisting of functions that have $k$ weak derivatives that are all in $L^{p}([a_{j},b_{j}])$ with respect to standard Lebesgue measure. In the special case where $\alpha_{j},\beta_{j}\in\{1/2,-1/2\}$ we allow $h_{j}\in W^{k,q}([a_{j},b_{j}])$, $k\geq 1$, $q>4$. Below, we reserve the index $k$ to refer to this assumed smoothness class.
As noted above, [40] considered the case of a single interval with $\alpha_{1},\beta_{1}\in\{1/2,-1/2\}$. It was shown that the error term is, in particular, $\mathrm{O}\left(\frac{\log n}{n^{k+\alpha}}\right)$ if $h_{1}\in C^{k,\alpha}([a_{1},b_{1}])$ and $k\geq 3$. Here we obtain similar error bounds [^1] In a previous version of this manuscript, an error bound without the $\log n$ factor was stated. While we believe this bound to still be valid, there was a flaw in the argument. while allowing $g>0$, $P>0$, and only requiring $k\geq 1$, see Theorem 3.1. There is a strong reason to believe this is still less than optimal, see Figures 4 and 5. These plots seem to indicate that this bound is suboptimal by a factor of $1/n$.
In the same vein as [40], the method described in this paper works by constructing close-to-analytic extensions of each function $h_{j}$ using polynomial approximation. Then, again borrowing directly from [40], we reuse an argument based on the Bernstein–Walsh inequality. Here it requires a bit more care in its implementation using a comparison, see Proposition A.1.
@@
The outline of the paper is as follows. In the remainder of this section, we introduce the Fokas–Its-Kitaev Riemann–Hilbert problem and fix more notation. In Section 2 we present the $\overline{\partial}$ deformations for the general Jacobi-type case $\alpha_{j},\beta_{j}>-1$. This leads to Theorem 2.6. In Section 3 we first present an improvement for the Chebyshev-like case $\alpha_{j},\beta_{j}\in\{-1/2,1/2\}$ in Theorem 3.1 and then discuss the addition of a finite number of point masses. Section 4 discusses the implied asymptotics for recurrence coefficients and uses this to present numerical experiments demonstrating that Theorem 3.1 is likely suboptimal. We also include five appendices that include technical developments.
@@
1.1. The asymptotics of the recurrence coefficients
We briefly review the implications of the asymptotic estimates obtained here by reviewing the formulae given in [18]. For error terms $E_{1},E_{2},E_{3}$, the recurrence coefficients for the orthonormal polynomials with respect to a measure $\mu$ (1.1) satisfy
$b_{n}(\mu)^{2}$ $=\frac{1}{\mathfrak{c}^{2}}\frac{\displaystyle\frac{\Theta_{2}(\infty;\mathbf{d}_{2};(n+1-P)\mathbf{\Delta}+\boldsymbol{\zeta})}{\Theta_{1}(\infty;\mathbf{d}_{2};(n+1-P)\mathbf{\Delta}+\boldsymbol{\zeta})}+E_{1}}{\displaystyle\frac{\Theta_{2}(\infty;\mathbf{d}_{2};(n-P)\mathbf{\Delta}+\boldsymbol{\zeta})}{\Theta_{1}(\infty;\mathbf{d}_{2};(n-P)\mathbf{\Delta}+\boldsymbol{\zeta})}+E_{2}},$
$a_{n}(\mu)$ $=\frac{\Theta_{1}^{(1)}(\mathbf{d}_{2};(n-P)\mathbf{\Delta}+\boldsymbol{\zeta})}{\Theta_{1}(\infty;\mathbf{d}_{2};(n-P)\mathbf{\Delta}+\boldsymbol{\zeta})}-\frac{\Theta_{1}^{(1)}(\mathbf{d}_{2};(n+1-P)\mathbf{\Delta}+\boldsymbol{\zeta})}{\Theta_{1}(\infty;\mathbf{d}_{2};(n+1-P)\mathbf{\Delta}+\boldsymbol{\zeta})}+\mathfrak{g}_{1}+E_{3}.$
For the definition of $\Theta_{j}$ see (D.2), and the superscript (1) denotes the $\mathrm{O}(1/z)$ term in the expansion of the function at $\infty$. The vectors $\mathbf{d}_{j}$, $\mathbf{\Delta}$, and $\boldsymbol{\zeta}$ are determined in (2.3) and the constant $\mathfrak{g}_{1}$ is the $\mathrm{O}(1/z)$ term in the expansion of $\mathfrak{g}$ at $\infty$ (see (A.2)). The size of the error terms $E_{1},E_{2},E_{3}$ is determined by the size of the error terms, with respect to $n$, obtained in Theorem 2.6 or Theorem 3.1.
@@ 1.2. The Fokas–Its–Kitaev Riemann–Hilbert problem
(1.6) $\mathrm{Res}_{z=c_{j}}\mathbf{Y}_{n}(z;\mu)$ $=\lim_{z\to c_{j}}\mathbf{Y}_{n}(z;\mu)\begin{bmatrix}0&\frac{{r}_{j}}{2\pi\mathrm{i}}\\ 0&0\end{bmatrix},\quad j=1,2,\ldots,P.$
@@
(1.8) $\begin{split}\mathbf{Y}_{n}(z;\mu)&=\begin{cases}\mathrm{O}\begin{bmatrix}1&1+|z-b_{j}|^{\alpha_{j}}\\ 1&1+|z-b_{j}|^{\alpha_{j}}\end{bmatrix}&\alpha_{j}\neq 0,\vskip 4.0pt\\ \mathrm{O}\begin{bmatrix}1&\log|z-b_{j}|\\ 1&\log|z-b_{j}|\end{bmatrix}&\alpha_{j}=0,\end{cases}\quad z\to b_{j},\\ \mathbf{Y}_{n}(z;\mu)&=\begin{cases}\mathrm{O}\begin{bmatrix}1&1+|z-a_{j}|^{\beta_{j}}\\ 1&1+|z-a_{j}|^{\beta_{j}}\end{bmatrix}&\beta_{j}\neq 0,\vskip 4.0pt\\ \mathrm{O}\begin{bmatrix}1&\log|z-a_{j}|\\ 1&\log|z-a_{j}|\end{bmatrix}&\beta_{j}=0,\end{cases}\quad z\to a_{j}.\end{split}$
@@
will be analytic except for possibly $a_{j},b_{j}$ where isolated singularities could persist. Indeed, we check, as $z\to b_{j}$, if $\alpha_{j}\neq 0$, for example,
@@
Remark 1.2.
We typically use $f^{\pm}(z)$ to denote the boundary values of a function $f$ from the left (++) or right (-) side of an oriented curve. In some cases, it may be convenient to use subscripts to denote the boundary values, e.g., $f^{\prime}_{\pm}(z)$.
@@ 2. Dbar deformation
We begin the deformation procedure by first recalling the definition and properties of the so-called $g$-function, see (A.2). Historically speaking, this construction is found in [2]. This function is also related to the exterior Green’s function with pole at $\infty$ [38]. Often this function is determined during the process of deformation, but we need it at the outset. Throughout what follows, we now suppose that $\epsilon^{\prime}>\epsilon>0$ are sufficiently small so that both $\varphi(z;a_{j}),\varphi(z;b_{j})$ from (A.4) are conformal on the balls $B_{\epsilon^{\prime}}(a_{j}),B_{\epsilon^{\prime}}(b_{j})$ respectively [^3] Here we use the notation $B_{\epsilon}(c)=\{z\in\mathbb{C}~:~|z-c|<\epsilon\}$, for all $j$. We assume that all the sets $B_{\epsilon^{\prime}}(a_{j}),B_{\epsilon^{\prime}}(b_{j})$, $j=1,2,\ldots,g+1$ are at a distance at least, say, $10\epsilon^{\prime}$ from one another. Next, define
@@
$\Gamma_{1}$$\Gamma_{3}$$\Gamma_{2}$$\sigma$0
@@
$\Sigma_{3,a_{j}}$$\Sigma_{1,a_{j}}$$\Sigma_{2,a_{j}}$$a_{j}$$\theta_{2}$$\theta_{1}$
$\Sigma_{1,b_{j}}$$\Sigma_{3,b_{j}}$$\Sigma_{2,b_{j}}$$b_{j}$
Figure 2. Mapping the contours $\Gamma_{j}$ using the local conformal mappings, with valid choices for $\theta_{1},\theta_{2}$ displayed.
The contours in Figure 2 will become jump contours in a deformed hybrid RH-$\overline{\partial}$ problem.
@@
In Appendix B, we develop a operator that extends a function off the real axis, while satisfying several convenient conditions. For $f_{j}(x)=1/h_{j}(x)$, consider $f_{j}^{\sharp}(x,y)$ with $\kappa=k$, $\tau={\epsilon/2}$. The choice of $\theta_{1},\theta_{2}$ that appear in the definition of $f_{j}^{\sharp}$ will be discussed below.
@@
giving a contour $\Sigma_{j,-}$, using left-to-right orientation. And a straight line, lying within the upper-half plane, can be used to connect
@@
giving a contour $\Sigma_{j,+}$, using left-to-right orientation. This is depicted in Figure 3. The contours $\Sigma_{j,\pm}$ now enclose open regions $\Omega_{j,\pm}$, also depicted in Figure 3.
$\Sigma_{3,a_{j}}$$\Sigma_{1,a_{j}}$$\Sigma_{2,a_{j}}$$a_{j}$$\Sigma_{1,b_{j}}$$\Sigma_{3,b_{j}}$$\Sigma_{2,b_{j}}$$b_{j}$$\Sigma_{j,+}$$\Sigma_{j,-}$$\Omega_{j,+}$$\Omega_{j,-}$
@@
$\overline{\partial}=\frac{1}{2}(\partial_{x}+\mathrm{i}\partial_{y}).$
@@
    For $c=a_{j},b_{j}$, $|z-c|<\epsilon/2$, and $z\in\Sigma_{c,1}\cup\Sigma_{c,3}$, we have f_j^♯(x,y) = f_j(z;c) = ∑_ℓ=0^k fj(ℓ)(c)ℓ! (z - c)^ℓ,
@@
    For $(x,y)\in{\Omega_{j,+}\cup\Omega_{j,-}}$, $|\overline{\partial}f_{j}^{\sharp}(x,y)|\leq C\|f_{j}\|_{C^{k,1}([a_{j},b_{j})]}|y|^{k}$, for some $C>0$.
We further suppose that $\epsilon$ is sufficiently small so that $f_{j}(z;c)$ does not vanish for $|z-c|\leq\epsilon/2$.